Evaluate f'(x)=for f(x)=x^2-2x at x=3

Found 2 solutions by Theo, Alan3354:
Answer by Theo(13342)

equation is f(x) = x^2 - 2x
when x = 3, you replace x with 3 to get:
f(3) = 3^2 = 2*3 = 9 - 6 = 3

Answer by Alan3354(69443)

Evaluate f'(x)=for f(x)=x^2-2x at x=3
------------------
f(x)=x^2-2x
f'(x) = 2x - 2
f'(3) = 2*3 - 2 = 4